# Reinforcement Learning from Bagged Reward

### v2

@@ Appendix B Omitted Proof
Consider a BRMDP where the bagged reward is defined as the sum of rewards for state-action pairs from the corresponding MDP contained within the same bag. In this context, the set of optimal policies for the standard MDP $\Pi$, aligns with that of the BRMDP $\Pi_{\mathrm{B}}$, implying that $\Pi=\Pi_{\mathrm{B}}$.
@@
Let us consider the structure of bagged rewards within BRMDP. For a given bag $\mathit{{B}_{i,n_{i}}}$, the bagged reward $R(\mathit{{B}_{i,n_{i}}})$ is the sum of the individual rewards of the state-action pairs contained within it, given by:
$R(\mathit{{B}_{i,n_{i}}})=\sum_{t=i}^{i+n_{i}-1}r(s_{t},a_{t}).$
@@
$\sum_{B\in\mathcal{B}_{\tau}}R(B)=\sum_{i\in\mathcal{I}_{\tau}}R(\mathit{{B}_{i,n_{i}}})=\sum_{i\in\mathcal{I}_{\tau}}\bigg{(}\sum_{t=i}^{i+n_{i}-1}r(s_{t},a_{t})\bigg{)}=\sum_{t=0}^{T-1}r(s_{t},a_{t}),$
@@
The policy optimization objective in BRMDP, $J_{\mathrm{B}}(\pi)$, aims to maximize the expected sum of bagged rewards along the trajectory from $t=0$, which is equivalent to maximizing the standard cumulative reward in MDP, $J(\pi)$:
@@
$=\mathbb{E}_{\pi,P,\mu}\bigg{[}\sum_{t=0}^{T-1}r(s_{t},a_{t})\bigg{]}=J(\pi).$
Given that the expected cumulative rewards for any policy in the BRMDP and MDP frameworks are equivalent, under the condition of infinite exploration or exhaustive sampling within the state-action space, the sets of optimal policies for each framework also coincide, implying that $\Pi=\Pi_{\mathrm{B}}$. ∎
Appendix C Experiment Settings and Implementation Details
Benchmarks with Bagged Rewards.
In this setting, the agent received a bagged reward at the end of each collected bag. Reward bag experiments of different bag sizes (5, 25, 50, 100, 200, and 500) and trajectory feedback were set up to verify the effectiveness of the method. To evaluate the efficacy of proposed method, commonly used trajectory feedback algorithms were adapted to fit the bagged reward setting as baselines. In these experiments, each reward bag was treated as an individual trajectory, and these modified algorithms were applied accordingly. Additionally, experiments using standard trajectory feedback were conducted to provide a comparative baseline within the unique setting. The total episodic feedback was computed at the end of the trajectory and was the sum of the per-step rewards the agent had collected throughout the episode. This experiment setting was the same as some previous works for learning from trajectory feedback Gangwani et al. (2020); Ren et al. (2021).
Implementation Details and Hyper-parameter Configuration.
@@ Appendix C Experiment Settings and Implementation Details
Appendix D Additional Experimental Results
This section provides further analysis and insights through additional experiments to complement the main findings presented in the paper.
Figure 6: Performance comparison adding SAC with reward shaping (SAC-shaping) in the Hopper-v2 and Walker2d-v2 environments, displaying both the mean and standard deviation over 6 trials with different random seeds. The experiments are conducted across six specific fixed bag lengths: 5, 25, 50, 100, 200, and 500, as well as with trajectory feedback, labeled as $9$, in each environment.
@@ Appendix D Additional Experimental Results
D.1 Experimental Result of Various-Length Reward Bags
In Table 3, we present the experiment results on various-length reward bags. The experiment depict in the table showcases the results of various methods applied across different environments with varying bag lengths of rewards, where bags are one next to another as in the definition of RLBR. This experiment reveals that longer bags tend to degrade the performance of most methods. However, our RBT method appears to be less sensitive to changes in bag length, maintaining robust performance even when the bag length is equal to the full trajectory. This result aligns with the result in Section 5.1.
D.2 Ablation Study
D.3 Rewards Comparison in Hopper-v2 with Bagged Rewards
@@
D.4 Reward Shaping
To complement our findings, we conducted additional experiments on reward shaping, following the naive shaping approach described in Hu et al. (2020). These experiments were performed in the Hopper-v2 and Walker2d-v2 environments due to the availability of task-specific weights from the original paper. As shown in the Fig. 6, the performance of SAC with reward shaping (SAC-shaping) varies with different bag lengths.
At shorter bag lengths, SAC-shaping tends to perform worse than standard SAC, possibly because the more frequent but less informative rewards add noise to the training process. However, as bag length increases, SAC-shaping slightly outperforms SAC, likely because the shaped rewards offer clearer, longer-term signals that assist in learning effective policies when rewards are less frequent. Essentially, reward shaping at longer bag lengths creates a more stable and beneficial learning signal for policy improvement. The figure demonstrates that while reward shaping improves SAC performance with larger bag sizes, our RBT method consistently outperforms both. RBT excels because it not only smooths the rewards but also redistributes them by considering the context within each bag. This comprehensive approach allows RBT to more accurately capture environmental complexities, leading to better performance across different bag lengths, indicating its robustness and efficiency in utilizing extended sequences of data.
D.5 Architecture Sensitivity

### v3

@@ Appendix B Omitted Proof
Provide a BRMDP where the bagged reward is regarded as the sum of rewards for state-action pairs of a MDP after reward redistribution contained within the same bag. In this context, the set of optimal policies for the BRMDP $\Pi_{\mathrm{B}}$ aligns with that of the MDP after reward redistribution $\Pi$, implying that $\Pi=\Pi_{\mathrm{B}}$.
@@
For a given bag $\mathit{{B}_{i,n_{i}}}$, if the bagged reward $R(\mathit{{B}_{i,n_{i}}})$ is regarded as the sum of the individual rewards of the state-action pairs contained within it, we can get:
$R(\mathit{{B}_{i,n_{i}}})=\sum_{t=i}^{i+n_{i}-1}\hat{r}(s_{t},a_{t}).$
@@
$\sum_{B\in\mathcal{B}_{\tau}}R(B)$ $=\sum_{i\in\mathcal{I}_{\tau}}R(\mathit{{B}_{i,n_{i}}})$
$=\sum_{i\in\mathcal{I}_{\tau}}\bigg{(}\sum_{t=i}^{i+n_{i}-1}\hat{r}(s_{t},a_{t})\bigg{)}=\sum_{t=0}^{T-1}\hat{r}(s_{t},a_{t}),$
@@
The policy optimization objective in BRMDP, $J_{\mathrm{B}}(\pi)$, aims to maximize the expected sum of bagged rewards along the trajectory from $t=0$, which is equivalent to maximizing the cumulative reward in redistributed reward MDP, $J(\pi)$:
@@
$=\mathbb{E}_{\pi,P,\mu}\bigg{[}\sum_{t=0}^{T-1}\hat{r}(s_{t},a_{t})\bigg{]}=J(\pi).$
Given that the expected cumulative rewards for any policy in the original BRMDP and redistributed reward MDP frameworks are equivalent, under the condition of infinite exploration or exhaustive sampling within the state-action space, the sets of optimal policies for each framework also coincide, implying that $\Pi=\Pi_{\mathrm{B}}$. ∎
@@
Appendix C Experiment Settings and Implementation Details
Benchmarks with Bagged Rewards.
Reward bag experiments of different bag sizes (5, 25, 50, 100, 200, and 500) and trajectory feedback were set up to verify the effectiveness of the method. To evaluate the efficacy of proposed method, commonly used trajectory feedback algorithms were adapted to fit the bagged reward setting as baselines. In these experiments, each reward bag was treated as an individual trajectory, and these modified algorithms were applied accordingly. Additionally, experiments using standard trajectory feedback were conducted to provide a comparative baseline within the unique setting. The total episodic feedback was computed at the end of the trajectory and was the sum of the per-step rewards the agent had collected throughout the episode. This experiment setting was the same as some previous works for learning from trajectory feedback Gangwani et al. (2020); Ren et al. (2021).
For the sparse reward experiments, the ball_in_cup-catch task provides a sparse reward of 1 when the ball is in the cup and 0 otherwise; in the reacher-hard task, the reward is 1 when the end effector penetrates the target sphere Tassa et al. (2018). Therefore, we treated the sparse reward as trajectory feedback and conducted experiments where the entire trajectory was considered as a single reward bag in these environments. The bagged reward was identical to the sparse reward provided by the environment.
Implementation Details and Hyper-parameter Configuration.
@@ Appendix C Experiment Settings and Implementation Details
Figure 6: Comparison of proposed method and sparse reward baselines. The mean and standard deviation are displayed over 6 trials with different random seeds across a total of 1e6 time steps.
@@
Appendix D Discussion on Sparse Reward
The sparse reward setting presents significant challenges due to infrequent feedback, making it difficult for agents to effectively explore the environment and discover successful strategies. To address this challenge, various methods have been developed to enhance exploration. Reward shaping strategies Ng et al. (1999); Hu et al. (2020); Tambwekar et al. (2019) added rewards to actions in a way that guides the agent towards better policies without altering the original reward function. Curiosity-driven methods Pathak et al. (2017); Sekar et al. (2020) encouraged agents to explore the environment by visiting unseen states, potentially solving tasks with sparse rewards. Additionally, curriculum learning in RL Florensa et al. (2018); Riedmiller et al. (2018) involved presenting an agent with a sequence of tasks with gradually increasing complexity, allowing the agent to eventually solve the initially given sparse reward task.
Due to the formal similarity between sparse rewards and trajectory feedback, both involving a single reward signal after a sequence, these methods can be directly applied to RLTF. However, the key difference lies in the nature of the rewards: sparse rewards typically focus on specific key points, making the reward Markovian, whereas RLTF evaluates the entire trajectory, indicating a Non-Markovian reward. Furthermore, sparse reward strategies address the exploration-exploitation trade-off inherent in sparse rewards, while RLTF focuses on distributing the evaluation of the entire trajectory across specific state-action pairs to guide learning. Thus, while these methods are applicable, they do not address the unique challenges posed by bagged rewards, as they fail to account for the reward structure within the bags. Therefore, they are not ideal solutions for RLTF Gangwani et al. (2020); Ren et al. (2021).
In Fig. 6, we present additional results comparing our method with sparse reward baselines. Shaping Tessler et al. (2019); Hu et al. (2020) and LIRPG Zheng et al. (2018) are as described in the main paper. The Random Network Distillation (RND) Burda et al. (2019) method introduces an exploration bonus for deep reinforcement learning by measuring the prediction error of a randomly initialized neural network, which is a widely recognized and utilized approach for addressing sparse reward challenges. We include it as a baseline in our comparison of sparse reward methods for a more comprehensive evaluation.
These results demonstrate that, despite the superficial similarity between sparse rewards and bagged rewards, the internal structure of the rewards differs, leading to suboptimal performance of sparse reward baselines in bagged reward environments. Conversely, as shown in the Section Compare with SOTA Methods of the main paper, our method performs well even in sparse reward environments, surpassing baselines specifically designed for sparse reward scenarios. This highlights the broad applicability and adaptability of our approach across different reward settings.
Appendix E Additional Experimental Results
This section provides further analysis and insights through additional experiments to complement the main findings presented in our main paper.
E.1 Experimental Result of Various-Length Reward Bags
In Table 4, we present the experiment results on various-length reward bags. The experiment depict in the table showcases the results of various methods applied across different environments with varying bag lengths of rewards, where bags are one next to another as in the definition of RLBR. This experiment reveals that longer bags tend to degrade the performance of most methods. However, our RBT method appears to be less sensitive to changes in bag length, maintaining robust performance even when the bag length is equal to the full trajectory. This result aligns with the result in the Section Compare with SOTA Methods in our main paper.
E.2 Rewards Comparison in Hopper-v2 with Bagged Rewards
@@
E.3 Architecture Sensitivity